# Resolving the Ambiguity of Complete-to-Partial Point Cloud Registration for Image-Guided Liver Surgery with Patches-to-Partial Matching

### v1

@@ V-B In Vitro Phantom Validation
TABLE III: Comparison of registration errors on the In Vitro phantom dataset across different visibility ratios. The average RMS-TRE ± standard deviation is reported in millimeters.
@@
Figure 6: Qualitative comparison of registration results on the in vitro phantom dataset. The source point cloud and its fiducial markers are shown in blue, while the target point cloud and its fiducial markers are shown in red, with fiducial markers displayed in larger sizes for visibility.
@@ V-B1 Performance against Baselines
We further compare the baseline methods with and without our proposed module using the phantom datasets. The results are reported in Table III. As seen, the In Vitro results align with the findings from the In Silico experiments. GO-ICP, RegTr, and RoITr are more sensitive to visibility variations compared to LiverMatch and Lepard. After applying our module, the RMS-TRE values for LiverMatch and Lepard are reduced from $20.54$ mm to $14.97$ mm and from $20.71$ mm to $12.45$ mm, respectively. These improvements are statistically significant.
@@ V-C Sensitivity and Ablation study
Figure 7: Sensitivity study on cluster number K, performed on the in vitro phantom dataset at different visibility ratio, using Lepard as the baseline for the proposed module.
TABLE IV: Ablation study on the selection method and candidate proposal, performed on the in vitro phantom dataset at a visibility range of [0.2, 0.3), using Lepard as the baseline for the proposed module. The average RMS-TRE ± standard deviation is reported in millimeters.
Selection rule Inlier number 14.14 $\pm$ 16.70
Closest point distance (default) 12.45 $\pm$ 15.72
Candidate proposal w/o 14.91 $\pm$ 19.06
w/ (default) 12.45 $\pm$ 15.72
We further conduct comprehensive sensitivity and ablation studies on the phantom dataset, incorporating the proposed module with Lepard to evaluate its effectiveness.
@@
Selection role. We further compare our closest distance-based selection introduced in Eq. 3, with the inlier number-based selection in Eq. 2, where the parameter $\tau$ is set to 0.05. As shown in Table IV, our approach proves to be more effective, resulting in fewer errors below 2 mm.
Patch candidate proposal. We also investigate the contribution of the candidate proposal to our proposed module. As shown in Table IV, omitting the patch candidate proposal results in a significant decline in performance, with a drop of approximately 2.5 mm in registration accuracy.
@@
Using the correspondence-based RANSAC ICP implementation from Open3D [43], with the hyperparameter for maximum correspondence distance set to 0.05, we found that our proposed module consistently outperforms RANSAC, particularly in low-visibility scenarios. RANSAC slightly worsened registration results in the visibility range $[0.4,1)$, likely due to its random sampling nature. The average running time of the proposed module is 0.04 seconds, compared to 1.06 seconds for RANSAC.
VI Discussion
@@
To alleviate this ambiguity, we propose a patches-to-partial feature matching module to bridge the gap between methods and their applications in liver surgery. Also, as a comprehensive benchmark is missing in this field, we construct a benchmark consisting of in silico and in vitro datasets.
We then evaluate representative correspondence-based methods on the constructed datasets. While these methods achieve comparable registration errors in high-visibility cases, their performance deteriorates as the visibility ratio decreases. Among them, RegTr is particularly sensitive to reduced visibility, likely due to its unique design. RegTr directly predicts the displacements of source and target key points as correspondences, rather than matching features between the two point clouds. This approach works well when the source and target point clouds are of similar scale, but in cases of complete-to-partial ambiguity, predicting key point displacements is less robust due to the lack of global context from the target point cloud. RoITr, a method based on the node-to-group strategy, also shows sensitivity to low-visibility cases. Its node-to-group strategy is not designed to handle complete-to-partial ambiguity but rather aims to find more correspondences through a coarse-to-fine search. In comparison with RegTr and RoITr, LiverMatch and Lepard are less sensitive to decreasing visibility ratios. Both methods employ a U-Net-like structure and match features between the source and target point clouds at the final output, indicating that this simple pipeline may be more versatile. Despite showing promising results from LiverMatch and Lepard, their performances are still not satisfying in low-visibility cases. In such instances, after applying our module, we observe significant improvements $w.r.t.$ the registration performance (see Table I, III), having merits in:
@@
i) Our proposed module alleviates the complete-to-partial ambiguity by converting it to patches-to-partial matching. Our module only has one hyperparameter, $K$, whose physical meaning controls the number of proposed patches. From the parameter-sensitivity study shown in Fig. 7, increasing $K$ will improve the registration by expanding a fully parallelizable search space, and the default value $K=5$ yields satisfactory results. In addition, we conducted an ablation study on our candidate proposal, and the closest point distance-based selection rule effectively contributes to the whole strategy.
@@
Limitations. The registration results depend on the condition that the source and target point clouds have similar densities. This may not hold if a very sparse target point cloud is manually collected via optical trackers. However, point cloud resampling methods may alleviate this limitation.
Experiments conducted in this paper are based on our constructed in silico and in vitro datasets. While our experiments have promisingly established a solid foundation for real-world applications, in future work, we will seek possibilities to construct an in vivo dataset to validate the performance of correspondence-based methods further.
@@ VII Conclusion
We have constructed a benchmark using both in silico and in vitro datasets to evaluate state-of-the-art learning-based point cloud matching methods. Our analysis underscores the challenge of complete-to-partial ambiguity in scenarios with low visibility ratios, as typically occurs during liver surgery. In light of this view, we introduce a patches-to-partial strategy as a module that can be seamlessly integrated with learning-based point cloud correspondence methods.
We have demonstrated the effectiveness of this module in improving several methods, especially in low-visibility conditions. To summarize, the constructed benchmark and the proposed module lay a solid foundation for future studies of in vivo applications.

### v2

@@ V-B In Vitro Phantom Validation
TABLE VII: Comparison of registration errors on the in vitro phantom dataset across different visibility ratios. The average RMS-TRE ± standard deviation is reported in millimeters.
@@
Figure 10: Qualitative comparison of registration results on the in vitro phantom dataset. The source point cloud and its volumetric fiducial markers are shown in blue, while the target point cloud and its volumetric fiducial markers are shown in red, with fiducial markers displayed in larger sizes for visibility.
@@ V-B1 Performance against Baselines
We further compare the baseline methods with and without our proposed module using the phantom datasets. The results are reported in Table VII. As seen, the In Vitro results align with the findings from the In Silico experiments. GO-ICP, RegTr, and RoITr are more sensitive to visibility variations compared to LiverMatch and Lepard. After applying our module, the RMS-TRE values for LiverMatch and Lepard are reduced from $20.54$ mm to $14.97$ mm and from $20.71$ mm to $12.45$ mm, respectively. These improvements are statistically significant.
@@ V-C Sensitivity and Ablation study
We further conduct comprehensive sensitivity and ablation studies on the in vitro phantom dataset, incorporating the proposed module with Lepard to evaluate its effectiveness.
@@
Selection role. We further compare our closest distance-based selection introduced in Eq. 4, with the inlier number-based selection in Eq. 3, where the parameter $\tau$ is set to 0.05. As shown in Table V, our approach proves to be more effective, resulting in fewer errors below 2 mm.
Patch candidate proposal. We also investigate the contribution of the candidate proposal to our proposed module. As shown in Table V, omitting the patch candidate proposal results in a significant decline in performance, with a drop of approximately 2.5 mm in registration accuracy.
@@
Using the correspondence-based RANSAC ICP implementation from Open3D [45], with the hyperparameter for maximum correspondence distance set to 0.05, we found that our proposed module outperforms RANSAC in low-visibility scenarios, achieving a RMS-TRE that is 3.10 mm lower than RANSAC’s result.
RANSAC slightly worsened registration results in the visibility range $[0.4,1)$, likely due to its random sampling nature.
Running time. The average running time of the proposed module under the default setting ($K=5$) on the in vitro dataset with the visibility [0.2, 0.3) is 0.04 seconds, compared to 1.06 seconds for RANSAC.
Computational complexity. The approximated computational complexity of our P2P strategy is $\mathcal{O}(K\cdot M\cdot M)$, whereas a complete-to-partial registration has $\mathcal{O}(N\cdot M)$. The complexity of P2P increases with $K$, and both approaches have similar computational complexity when $K\cdot M=N$.
VI Discussions
VI-A Limitations of Baselines and Merits of P2P
@@
To alleviate this ambiguity, we propose a P2P feature matching module to bridge the gap between methods and their applications in liver surgery. Also, as a comprehensive benchmark is missing in this field, we construct a benchmark consisting of in silico and in vitro datasets.
We then evaluate representative correspondence-based methods on the constructed datasets. While these methods achieve comparable registration errors in high-visibility cases, their performance deteriorates as the visibility ratio decreases.
VI-A1 Limitation of End-to-End Correspondence Design
RegTr is particularly sensitive to reduced visibility, likely due to its unique design. RegTr directly predicts the displacements of source and target key points as correspondences, rather than matching features between the two point clouds. This approach works well when the source and target point clouds are of similar scale, but in cases of complete-to-partial ambiguity, predicting key point displacements is less robust due to the lack of global context from the target point cloud.
VI-A2 Limitation of Node-to-Group Strategy
RoITr, a method based on the node-to-group strategy [14, 27], also shows sensitivity to low-visibility cases. This strategy does not consider the complete-to-partial ambiguity and extracts patches from both the source and target point clouds, with a hyperparameter controlling patch size. In contrast, our method represents a source point cloud with patches that contain the same number of points as the target. Also, its Superpoint Matching module to generate patch nodes at a coarse scale introduces misleading nodes due to coarse-scale errors and complete-to-partial ambiguity. In our approach, we first localize a region and then sample nodes from it, leading to more reliable patch selection. Moreover, we utilize a distance-based selection rule (Eq. 4) to select the final rigid transformation, which demonstrates greater robustness than the inlier-based selection rule (Eq. 3) in handling complete-to-partial ambiguity (see Table V).
VI-A3 Limitation of Complete-to-Partial Matching
In comparison with RegTr and RoITr, LiverMatch and Lepard are less sensitive to decreasing visibility ratios. Both methods employ a U-Net-like structure and match features between the source and target point clouds at the final output, indicating that this simple pipeline may be more versatile. Despite showing promising results from LiverMatch and Lepard, their performances are still not satisfying in low-visibility cases.
VI-A4 Merits of the P2P Module
Given the above discussion, the designs of RegTr and RoITr are not well-suited for low-visibility cases. To address the complete-to-partial ambiguity, we developed the P2P module and integrated it into LiverMatch and Lepard.
In such instances, after applying our module, we observe significant improvements $w.r.t.$ the registration performance (see key quantitative improvements in Table I, VII). For example, under low visibility range [0.2, 0.3), the approach improved registration errors for LiverMatch from 12.9 mm to 8.7 mm (-33%) for its in silico datasets and from 20.5 mm to 15.0 mm (-27%) for its in vitro datasets. Similarly, for Lepard, registration errors were improved from 9.5 mm to 6.7 mm (-29%) for its in silico datasets and from 20.7 mm to 12.5 mm (-40%) for its in vitro datasets.
Similar trends are seen under noise and deformation (Table II, III). The P2P module consistently reduced RMS-TREs, with greater gains at higher noise. Under both low and high deformation ranges, P2P-enhanced methods showed lower errors and reduced variability. For instance, LiverMatch improved by $4.0$ mm and $4.7$ mm, and Lepard by $2.7$ mm and $2.7$ mm, respectively, for their in silico and in vitro datasets.
In addition to the significant improvement in terms of registration error, the P2P has merits in:
@@ VI-A4 Merits of the P2P Module
i) Our proposed module alleviates complete-to-partial ambiguity by converting it to P2P matching. Our module only has one hyperparameter, $K$, whose physical meaning controls the number of proposed patches. From the parameter-sensitivity study shown in Fig. 8, increasing $K$ will improve the registration by expanding a fully parallelizable search space, and the default value $K=5$ yields satisfactory results. The choice of $K=5$ is theoretically justified by the lowest characteristic visibility ratio, which is approximately 20% [17]. Given this minimum visibility, partitioning the complete point cloud into five patches ensures that at least one patch is likely to contain a corresponding structure in the partial target, even under the most challenging visibility conditions.
In addition, we conducted an ablation study on our candidate proposal, and the closest point distance-based selection rule effectively contributes to the whole strategy.
@@
VI-B Potential Challenges for Correspondence-based Point Cloud Registration
Correspondence-based point cloud registration methods typically require that the source and target point clouds share similar properties, such as density and reconstruction accuracy. However, these prerequisites may be challenged in the context of liver surgery due to intraoperative constraints and variability in data acquisition.
One prerequisite is that the target point cloud represents a continuous liver surface region, allowing it to be voxelized to achieve a density similar to that of the source. However, this assumption may not hold when the target point cloud is manually collected using optical trackers or when other organs and surgical instruments partially occlude the intraoperative surface. To mitigate this issue, interpolation-based approaches, such as the grid-fitting method proposed by Collins et al.[50], or point cloud completion methods [51] may help reconstruct a dense and continuous surface representation.
Minimizing reconstruction errors in both the source and target point clouds is crucial for accurate 3D-3D registration. In the source point cloud, errors may arise from liver segmentation from preoperative scans and the inherent limitations of scan resolution. For the target point cloud, inaccuracies often stem from depth sensor limitations and constraints in 3D surface reconstruction techniques. Despite these challenges, continuous advancements in hardware (e.g., high-precision depth sensors, CT) and software (e.g., stereo depth estimation algorithms[52]) are steadily improving the feasibility of this approach for liver surgery.
VI-C Limitations of Constructed Datasets
In our current in silico dataset generation pipeline, the target point cloud is created by randomly cropping the full liver surface. However, in open and laparoscopic liver surgery, only the anterior liver surface is visible due to occlusions from other organs and surgical instruments. This discrepancy between synthetic and real-world visibility may hinder model generalizability. A cropping method that better simulates how a camera perceives the anterior liver surface during surgery could improve model performance when trained on the in vivo dataset and testing on real surgical data.
Our in silico and in vitro datasets primarily feature moderate deformations, similar to those in the Sparse Non-Rigid Registration Challenge dataset [47]. The deformation simulation pipeline [21] used to generate the in silico dataset randomly applies zero boundary conditions and external forces, occasionally producing cases dominated by rigid transformations. Following the rigid alignment step described in §IV-A1, we observed that the resulting deformations, quantified by RMS-TREs and visualized in Fig. 4, are unevenly distributed and predominantly moderate.
In future work, we aim to build a in vivo dataset to further validate the performance of correspondence-based registration methods in real surgical scenarios, bringing our research closer to clinical applicability. Insights gained from constructing the in vivo dataset, such as the characteristic deformation levels observed in laparoscopic surgery and the zero-displacement boundary conditions imposed by anatomical structures, will inform refinements to the in silico dataset, to improve training efficacy and generalization performance.
VI-D Limitations of the P2P module
To reduce redundant patches, the default setting of the P2P module generates patch nodes from the top candidate set, a design proven to be more effective than their selection from the entire set (see Table V). However, this approach may produce candidate patches that fail to include the correct anatomical region in extreme cases, which may result from the prerequisites outlined in §VI-E being unmet or due to extreme deformation. This issue may arise when the source and target features become highly inconsistent, leading to the failure of Visible Source Point Estimation in §III-B. In such cases, disabling our patch candidate proposals and instead applying FPS directly to the anterior surface to generate patch nodes while increasing their number is advisable.
Currently, the P2P module acts as an additional module to the networks. However, designing a network with a built-in P2P strategy may enhance efficiency and effectiveness. Exploring this approach is part of our future work.
VI-E Clinical Significance
The clinical significance of this work lies in translating learning-based point cloud registration techniques to streamline image-guided liver surgery workflows. Unlike conventional approaches that require manual interaction for initial rigid registration, learning-based methods have the potential to provide fully automatic and highly accurate alignment. This approach not only reduces operative time, but also facilitates initialization for subsequent non-rigid registration methods [1, 48, 47] and aids surgeons in precisely localizing subsurface structures for more accurate tumor resection. Furthermore, given the anatomical similarities between the liver and other organs, such as the prostate and kidney, the proposed techniques and module could extend their benefits to image-guided surgeries for these organs as well.
@@ VII Conclusion
We have constructed a benchmark using both in silico and in vitro datasets to evaluate state-of-the-art learning-based point cloud matching methods. Our analysis underscores the challenge of complete-to-partial ambiguity in scenarios with low visibility ratios, as typically occurs during liver surgery. In light of this view, we introduce a P2P strategy as a module that can be seamlessly integrated with learning-based point cloud correspondence methods. We have demonstrated the effectiveness of this module in improving several methods, especially in low-visibility conditions.
We further discussed the limitations of state-of-the-art methods, our constructed datasets, and the proposed module in the context of in vivo applications. To address these limitations, future work will focus on optimizing the cropping and deformation simulation in the in silico dataset, constructing dedicated in vivo datasets, exploring interpolation and point cloud completion techniques for improved intraoperative surface reconstruction, and developing a network that natively integrates the P2P strategy within its feature extraction pipeline.
To summarize, the constructed benchmark, the proposed module, and our comprehensive discussion lay a solid foundation for future studies of in vivo applications.